40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 184     Entry time: Wed May 14 02:15:15 2014
Author: Koji 
Type: Optics 
Category: Characterization 
Subject: FSR/TSM adjustment of the OMC cavity  

1. FSR was adjusted and measured with "the golden arches" technique again.

FSR = 264.8412 MHz +/- 1400Hz => Lcav = 1.13197 m. (nominal 1.132m)

2. Transverse mode spacings for the vertical and horizontal modes were measured.

TMS/FSR = 0.218144 (V) / 0.219748 (H)

This is almost perfect!

The 19th-order lower sideband hit the resonance. Next step is to glue some of the flat mirrors.

Attachment 1: Cav_scan_response_140503_Pitch.pdf  140 kB  Uploaded Wed May 14 03:17:11 2014  | Show | Hide all | Show all
Attachment 2: Cav_scan_response_140503_Yaw.pdf  204 kB  Uploaded Wed May 14 03:18:57 2014  | Hide | Hide all | Show all
Cav_scan_response_140503_Yaw.pdf
ELOG V3.1.3-